OpenStudy (sissyedgar):

The length of one side of a square is square-root 86cm. What is the area of the square? A. 7396 cm2 B. 43 cm2 C. 172 cm2 D. 86 cm2

OpenStudy (sissyedgar):

@Whitemonsterbunny17

OpenStudy (loser66):

yes

OpenStudy (anonymous):

\[s=\sqrt{86}\] \[A=s^{2}\] so the area would be \[(\sqrt{86})^{2}\]

OpenStudy (anonymous):

which is *not* B...

OpenStudy (anonymous):

no...\[(\sqrt{a})^{2}=a\]

OpenStudy (sissyedgar):

i actually ment 2 say times 4 but ok

OpenStudy (sissyedgar):

so its c?

OpenStudy (sissyedgar):

a simple yes or no sil do... i really have to finish this assighnment

OpenStudy (sissyedgar):

will not sil

OpenStudy (anonymous):

it is not c.

OpenStudy (sissyedgar):

im not good woth letter math.... so is it a or d?

OpenStudy (sissyedgar):

i ment with not woth

OpenStudy (sissyedgar):

so is it A. or D.

OpenStudy (sissyedgar):

whitemonsterbunny17!!!!!!!!!!!!!!!

OpenStudy (anonymous):

i'll help you understand it, but i won't give you the answer...

OpenStudy (sissyedgar):

my stuff is commin in more than once 4 sum reason

OpenStudy (sissyedgar):

huhhhhhhhggggg

whitemonsterbunny17 (whitemonsterbunny17):

Okay, okay. I don't know how to do this kind of math, so I can't help you, but.... @mtbender74 and @surjithayer please be a little bit more descriptive, and go thoroughly through the steps with her. And also, kindness is a key. Be kind, and be helpful. Bunnies is out!!! (and watching) *^o^* /) /) O( ^.^)o *:・゚✧.゚・*:・゚✧.゚・* C(")(") ☁☁☁☁☁☁☁☁☁☁☁☁☁☁

OpenStudy (sissyedgar):

lol

OpenStudy (sissyedgar):

no1 named @serjithayer was in here

OpenStudy (anonymous):

i'm always willing to take people through the steps, but if they don't understand what i said, i need to know where they are lost. if they want to know *how* to do the problem, i'll spend as long as it takes to get them to understand, but if they are simply looking for me to give them the answer, i won't... that being said, do you still want my help @sissyedgar

OpenStudy (sissyedgar):

yes i do

OpenStudy (anonymous):

ok...so first things first...do you understand what square roots are?

OpenStudy (sissyedgar):

yes

OpenStudy (anonymous):

so \[\sqrt{9}=3\] since (3)(3)=9

OpenStudy (anonymous):

so if we looked at \[(\sqrt{9})^{2}\]that would be \[(\sqrt{9})^{2}=(3)^{2}=9\] still with me?

OpenStudy (sissyedgar):

how do u type square root in the chat. kinda off subject i know but it is really cool

OpenStudy (sissyedgar):

yes i am

OpenStudy (anonymous):

you can use the equation button...or once you know how they are built, you can type them in directly :)

OpenStudy (anonymous):

so what we have is that the square of the square root of any number is just the number...in other words \[(\sqrt{a})^{2} = a\] regardless of the value of a

OpenStudy (sissyedgar):

ok i really need to go because my mother is saying i need to hurry up and finish so we can go!!!

OpenStudy (anonymous):

well...provided a is not negative...but don't worry about that

OpenStudy (anonymous):

what the problem said is that the side of the square is \[s=\sqrt{86}\]

OpenStudy (anonymous):

the formula for the area of a square is \[A=s^{2}\]

OpenStudy (sissyedgar):

srry i will come bk n get the answer but i have to go now

OpenStudy (anonymous):

so, the area of the square is \[A=(\sqrt{86})^{2}\]

OpenStudy (anonymous):

now, using the fact that \[(\sqrt{a})^{2}=a\], what do we have?